Let V be the vector space of polynomials of degree less than of equal to n.Show that the derivative operator on V is nilpotent of index n+1.


1
Expert's answer
2021-01-12T16:03:46-0500

Let P denotes a polynomial and D denotes the derivative operator. Then let "P=a_0 +a_1X+\cdots+a_kX^k" for "k\leq n." Then "DP=a_1+2a_2X+\cdots+ka_kX^{k-1}." Then "deg DP=k-1." Hence applying this repetitively we get, deg "DP=k-i" for "i\leq k" . For "i=k," the degree is 0 hence a constant polynomial. So for "i>k" the polynomial vanishes. Hence for any polynomial of degree "\leq n," "D^{n+1}(P)=0." Also for the ploynomial "X^{n}" , we see, "D^{n}(X^{n})=n!." Hence "D^{n}" is not the zero operator on V. Hence it is nilpotent of index n+1.